OpenStudy (anonymous):

(2x+1)(x-6) simplify by using distributed property

OpenStudy (mathstudent55):

The distributive property is: a(b + c) = ab + ac a(b - c) = ab - ac Here a = 2x + 1, and b = x - 6 (2x+1)(x-6) = (2x + 1) * x - (2x + 1) * (-6) = x(2x - 1) - (-6)(2x - 1) Now use the distributive property again for each set of perentheses.

OpenStudy (anonymous):

i dont get it

OpenStudy (mathstudent55):

Here's a simple example of the distributive property: 5(6 + 3) = 5 * 6 + 5 * 3 Do you understand so far?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Ok. Your problem is similar just a bit more complicated.

OpenStudy (mathstudent55):

This is the simple example again.